OpenStudy (anonymous):

Simplify. (k^7)^3 A) k^-10 B) k^4 C) k^10 D) k^21

OpenStudy (anonymous):

@Love_Ranaa

OpenStudy (jhannybean):

distributive property

OpenStudy (jhannybean):

\[\huge (x^n)^m = x^{m\cdot n}\]

OpenStudy (igreen):

Yep, multiply the exponent on the outside to all of the exponents on the inside.

OpenStudy (kendricklamar2014):

Mulitiply the Exponents: \[7*3 = ?\]

OpenStudy (anonymous):

21

OpenStudy (igreen):

Yes, so what's our final answer?

OpenStudy (kendricklamar2014):

Yes

OpenStudy (anonymous):

So D

OpenStudy (igreen):

Correct

OpenStudy (kendricklamar2014):

Yes

OpenStudy (jhannybean):

\(\checkmark\)

OpenStudy (anonymous):

Thank you guys!

OpenStudy (kendricklamar2014):

\(\color{blue}{You're ~Welcome}\)

OpenStudy (love_ranaa):

Sorry, I didn't see it